OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!! Consider the piecewise continuous function \[f(t)=\left\{\begin{matrix} t-3, &t<2 \\ 1, & t\geq 2 \end{matrix}\right.\] Rewrite f(t) in terms of unit step functions. Solve the IVP \[y''-4y'+4y=f(t), y(0)=-1, y'(0)=-2\]

OpenStudy (anonymous):

Make use of the Laplace transform. You may want to consider writing the given \(f(t)\) in terms of the Heaviside step function.

OpenStudy (anonymous):

^though doing so isn't absolutely necessary. \(f\) is simple enough to have each piece's transform computed quite easily from the integral definition.

OpenStudy (anonymous):

\[(u-t)U_{2}(t)+U_{2}(t)\] Is that how you re-write the piecewise

OpenStudy (anonymous):

I'm not sure I understand your notation. What's the distinction between \(u\) and \(U\)?

OpenStudy (anonymous):

I denote the step function as \[u(t-c)=\begin{cases}1&\text{for }t\ge c\\0&\text{for }t<c\end{cases}\] You should have \[f(t)=(t-3)u(-t+2)+u(t-2)\]

OpenStudy (anonymous):

for some reason that's how my professor showed us how to do but when i read the book, it's the same as you have it. so am a little bit confused

OpenStudy (anonymous):

Your prof might be using a slightly different notation, where \(u_c(t)=u(t-c)\) mean the same thing.

OpenStudy (anonymous):

Oh okay. that's now makes sense because it has been confusing me!

OpenStudy (anonymous):

\[f(t)=(t-3)u(-t+2)+u(t-2)\] \[\begin{align*} \mathcal{L}\{f(t)\}&=\int_0^\infty (t-3)u(-t+2)e^{-st}~dt+\int_0^\infty u(t-2)e^{-st}~dt\\\\ &=\int_0^2 (t-3)e^{-st}~dt+\int_2^\infty e^{-st}~dt\\\\ &=\left[-\frac{1}{s}(t-3)e^{-st}\right]_0^2+\frac{1}{s}\int_0^2 e^{-st}~dt+\int_2^\infty e^{-st}~dt\\\\ &=\frac{e^{-2s}-3}{s}+\frac{1-e^{-2s}}{s^2}+\frac{e^{-2s}}{s} \end{align*}\] and keep going from there.